OpenStudy (anonymous):

Solve and describe the steps you took to solve: 68=5m-7

OpenStudy (rina.r):

so, bring 7 to the right with 68 it will be +7

OpenStudy (rina.r):

so, 68+7=5m 75=5m m=75/5 =15

OpenStudy (unklerhaukus):

another way to describe the first step would be to add 7 to both sids of the equation

OpenStudy (anonymous):

Ah, ok. Thanks for the help!!